### v1

@@
Abstract
We consider a practically motivated variant of the canonical online fair allocation problem: a decision-maker has a budget of perishable resources to allocate over a fixed number of rounds. Each round sees a random number of arrivals, and the decision-maker must commit to an allocation for these individuals before moving on to the next round. The goal is to construct a sequence of allocations that is envy-free and efficient. Our work makes two important contributions toward this problem: we first derive strong lower bounds on the optimal envy-efficiency trade-off that demonstrate that a decision-maker is fundamentally limited in what she can hope to achieve relative to the no-perishing setting; we then design an algorithm achieving these lower bounds which takes as input $(i)$ a prediction of the perishing order, and $(ii)$ a desired bound on envy. Given the remaining budget in each period, the algorithm uses forecasts of future demand and perishing to adaptively choose one of two carefully constructed guardrail quantities. We demonstrate our algorithm’s strong numerical performance — and state-of-the-art, perishing-agnostic algorithms’ inefficacy — on simulations calibrated to a real-world dataset.
@@ 1 Introduction
Despite a consistent decline in food insecurity in the United States over the past decade, 2022 saw a marked increase in individuals struggling to access enough food to fulfill basic needs. A recent report by the U.S. Department of Agriculture found that over 44 million individuals faced some form of hunger in 2022 — 45% more than the previous year (Godoy, , 2023). Due in part to rising food prices and the rolling back of pandemic-era social security measures, this disturbing statistic has further underscored the important role of local food banks; for instance, the Feeding America network of food banks, food agencies, and local food programs distributed over 5.2 billion meals that same year (Feeding America, , 2022).
@@
In distributing food throughout their operating horizon, food banks have two competing objectives: distributing as much food as possible to communities in need, and ensuring equitable access to donations. This tension has attracted much attention in the operations literature, with recent work characterizing the fundamental trade-offs between fairness and overall utility in sequential allocation problems (Bertsimas et al., , 2011; Donahue and Kleinberg, , 2020; Lien et al., , 2014; Manshadi et al., , 2023; Sinclair et al., , 2022). Understanding such trade-offs in theory is useful, as they allow a system designer to recognize and choose their desired operating point, balancing the loss in efficiency and equity. Despite the useful insights derived from prior work, to the best of our knowledge an important reality of food bank operations remains overlooked: the existence of perishable goods, which constitute a substantial portion of food donations. The Los Angeles Regional Food Bank, for instance, distributed over 26 million pounds of produce in 2022 alone (Los Angeles Regional Food Bank, 2022a, ). Perishable goods pose significant challenges for these organizations, who frequently find themselves needing to throw out spoiled goods (Los Angeles Regional Food Bank, 2022b, ). Indeed, the equity-efficiency trade-off is exacerbated in the presence of perishables: while equity requires a decision-maker to allocate conservatively across arriving streams of demand (Sinclair et al., , 2022), perishability starts a “race against time.” As goods perish due to a slow allocation rate, not only is efficiency further harmed, but so may be equity, as spoilage runs the risk of a decision-maker running out of items, with nothing left to give out by the end of the operating horizon.
Thus motivated, this paper seeks to answer the following questions:
@@
Do established equity-efficiency trade-offs in dynamic environments persist in the presence of perishable goods? If not, what limits do they impose on fair and efficient allocations? Can we design policies that perform well under these limits?
Before detailing our contributions, we highlight that, though this work is motivated by food bank distribution systems, the interplay between fairness and perishability is an important consideration in several other settings, e.g., vaccine distribution (Manshadi et al., , 2023), electric vehicle charging (Gerding et al., , 2019), and federated cloud computing (Aristotle Cloud Federation Project, , 2022; Ghodsi et al., , 2011).
@@ 1.1 Our contributions
We consider a model in which a decision-maker has a fixed budget of items (also referred to as goods, or resources) to be allocated over $T$ discrete rounds. An a priori unknown number of individuals arrives in each period, each seeking a share of goods. Each agent is characterized by an observable type (drawn from a known, potentially time-varying distribution), associated with a linear utility function over the received allocation. Moreover, each unit of good has a stochastic perishing time (similarly drawn from a known distribution, independent of incoming arrivals), after which the good spoils and can no longer be allocated. The decision-maker allocates goods according to a fixed ordering (also referred to as perishing prediction, or allocation schedule), e.g., in increasing order of expected perishing time. The goal is to find a policy that trades off between three ex-post metrics:
@@
For this setting, we first characterize the fundamental limits of perishability placed on any online algorithm. We argue that — contrary to the setting without perishable resources — envy ceases to be a meaningful metric for a large class of perishing processes. To see this, consider an extreme scenario in which all items perish at the end of the first round. Clearly, there is no hope of achieving low envy in such a setting since future demand can never be satisfied. Our first main contribution identifies a necessary and sufficient condition on the joint perishing and arrival distribution for low counterfactual envy to be an achievable desideratum (Theorem 3.2). From a managerial perspective, this characterization — which at a high level states that the cumulative perishing must lag behind cumulative arrivals — can be viewed as guidance on the composition of perishables in the initial budget. It moreover underscores the importance of leveraging joint information over perishing and demand: if demand is back-loaded, perishing times should be concentrated late in the horizon; if most demand arrives early, however, early perishing times are acceptable.
@@
(a) Envy-Efficiency Pareto Frontier
[…]
(b) Dependence of Algorithm 1 performance on envy-perishing regime
Figure 1: Graphical representation of Theorems 4.2 and 3.7. Fig. 1(a) illustrates the envy-efficiency trade-off ($\Delta_{\text{\it efficiency}}$ vs. $\Delta_{\text{\it EF}}$) achieved by Perishing-Guardrail (Algorithm 1). The dotted lines represent the impossibility results due to either demand or perishing uncertainty. The region below the solid line represents the impossibility due to the envy-efficiency trade-off; the green region is the achievable region for Perishing-Guardrail. Fig. 1(b) illustrates the phase transition between the performance of Perishing-Guardrail depending on the spoilage loss $\mathcal{L}^{\textsf{perish}}$ ($x$-axis) and envy parameter $L_{T}$ ($y$-axis).
For this class of processes, which we term offset-expiring, zero spoilage occurs if the perishing prediction is perfect. We show, however, that inaccuracies in the allocation schedule pose an insurmountable barrier to any algorithm’s performance by characterizing an unavoidable loss in equity and efficiency due to these errors (Theorem 3.7). In contrast to the no-perishing setting, in which the only source of loss is exogenous uncertainty in demand, in our setting the loss due to spoilage is endogenous: it crucially depends on the rate at which the algorithm allocates items. This endogeneity poses significant challenges in the analysis of any adaptive algorithm; designing a tractable approach to analyzing ex-post spoilage is the main technical contribution of our work. Additionally, the lower bounds we derive give rise to the key insight that, contrary to the “race against time” intuition under which a decision-maker must increase the allocation rate to prevent avoidable spoilage, achieving low hindsight envy in the presence of unavoidable spoilage requires a decision-maker to potentially allocate significantly less than the proportional allocation. Hence, perishability throws a wrench into the well-studied envy-efficiency trade-off: while hindsight and counterfactual envy are aligned in the no-perishing setting, these two may be at odds when goods spoil, since only high counterfactual-envy solutions may yield low hindsight-envy. The tension between efficiency and equity is exacerbated for the same reason, relative to the classical setting.
In our final technical contribution, we leverage these insights to construct an adaptive threshold algorithm (Algorithm 1) that achieves these lower bounds (Theorem 4.2). Our algorithm takes as input $(i)$ the fixed allocation schedule, $(ii)$ a desired upper bound on hindsight envy $L_{T}$, and $(iii)$ a high-probability parameter $\delta$. Given these inputs, it computes a high-probability lower bound on a budget-respecting zero-hindsight-envy solution, and an “aggressive” efficiency-improving allocation that is $L_{T}$ away. In each round the algorithm chooses which of the two quantities to allocate to each individual, cautiously doing so by constructing pessimistic forecasts of future arrivals and perishing. While this algorithm is similar in flavor to state-of-the-art algorithms for the no-perishing setting (Sinclair et al., , 2022), the main challenge it contends with is forecasting (endogenous) future spoilage. Here, we leverage the bounding technique used to construct our lower bounds, which relies on the analysis of a knife-edge, “slow” consumption process that tractably decouples past allocations from future perishing. Our algorithm’s bounds give rise to three salient regimes (depicted graphically in Figure 1(b)), as a function of hindsight envy tolerance $L_{T}$ and the unavoidable loss due to spoilage per period, denoted by $\mathcal{L}^{\textsf{perish}}$:
@@
    Low-envy ($L_{T}\lesssim 1/\sqrt{T}$): there are no efficiency or counterfactual envy gains from deviating from the equitable solution $L_{T}=0$.
@@
    High-envy, high-perishing ($L_{T}\gtrsim 1/\sqrt{T},\mathcal{L}^{\textsf{perish}}\gtrsim 1/\sqrt{T}$): inefficiency is invariant to $L_{T}$; setting $L_{T}\sim\mathcal{L}^{\textsf{perish}}$ is optimal with respect to counterfactual envy.
@@
    High-envy, low-perishing ($L_{T}\gtrsim 1/\sqrt{T},\mathcal{L}^{\textsf{perish}}\lesssim 1/\sqrt{T}$): counterfactual envy increases as $L_{T}$, and inefficiency decreases as $1/L_{T}$, until reaching the unavoidable cumulative spoilage loss of $T\mathcal{L}^{\textsf{perish}}$.
These results further highlight the extent to which a decision-maker is restricted in leveraging inequity to improve efficiency (and vice versa).
We complement our theoretical bounds in Section 5 by testing the practical performance of our algorithm on both synthetic and real-world datasets. Our experiments show that the unfairness required to achieve these efficiency gains is order-wise larger than in settings without perishable resources. Additionally, they underscore the weakness of perishing-agnostic online algorithms. We observe that these latter algorithms are incapable of leveraging unfairness to improve efficiency across a variety of perishing regimes. In contrast to these, our algorithm’s construction of a perishing-aware baseline allocation $\underline{X}$ is necessary to mitigate stockouts across all — rather than simply worst-case — instances. These include instances where offset-expiry fails to hold with high probability, as is the case in the real-world dataset we use to calibrate our experiments (Keskin et al., , 2022). Perhaps most surprisingly, despite our baseline allocation being significantly lower than that of algorithms that don’t take into account perishability, our algorithm is more efficient than these more aggressive algorithms, in addition to being more fair. This observation contradicts the “race against time” intuition that aggressive allocations are necessarily more efficient than cautious ones. Finally, we numerically explore the question of setting practical allocation schedules that perform well along all metrics. Our main managerial insight is that ordering items in increasing order of a high-probability lower bound on their perishing time robustly trades off between the natural ordering that allocates items in increasing order of expected perishing time, and the inherent variability in the perishing process.
@@
We next survey the related literature. We present the model in Section 2, and formalize the limits of perishability in Section 3. In Section 4 we design and analyze an algorithm that achieves the derived lower bounds; we conclude with numerical experiments in Section 5.
@@ 1.2 Related work
Fairness in resource allocation has a long history in the economics and computation literature, beginning with Varian’s seminal work (Varian, , 1974, 1976). More recently, there has been ongoing work studying the intersection of fairness and operations, including assortment planning (Chen et al., , 2022), pricing (Cohen et al., , 2022; den Boer et al., , 2022), incentive design (Freund and Hssaine, , 2021), algorithmic hiring (Salem and Gupta, , 2023), and societal systems more generally (Gupta and Kamble, , 2021). We highlight the most closely related works below, especially as they relate to online fair allocation; see Aleksandrov and Walsh, 2019b for a survey.
@@
There exists a long line of work in which the non-perishable resource becomes available to the decision-maker online, whereas agents are fixed (Benade et al., , 2018; Aleksandrov et al., , 2015; Mattei et al., , 2017, 2018; Aleksandrov and Walsh, 2019a, ; Banerjee et al., , 2020; Bansal et al., , 2020; Bogomolnaia et al., , 2022; He et al., , 2019; Aziz et al., , 2016; Zeng and Psomas, , 2020). These models lie in contrast to the one we consider, wherein resources are fixed and individuals arrive online. Papers that consider this latter setting include Kalinowski et al., (2013), which considers maximizing utilitarian welfare with indivisible goods; Gerding et al., (2019) considers a scheduling setting wherein agents have fixed and known arrival and departure times, as well as demand for the resource; Hassanzadeh et al., (2023) allows individuals to arrive in multiple timesteps. A series of papers also consider the problem of fair division with minimal disruptions relative to previous allocations, as measured by a fairness ratio, a competitive ratio analog of counterfactual envy in our setting (Friedman et al., , 2017; Cole et al., , 2013; Friedman et al., , 2015). Other works design algorithms with attractive competitive ratios with respect to Nash Social Welfare (Azar et al., , 2010; Banerjee et al., , 2020), or the max-min objective (Lien et al., , 2014; Manshadi et al., , 2023).
@@
The above papers situate themselves within the adversarial, or worst-case, tradition. A separate line of work considers fair resource allocation in stochastic settings (Donahue and Kleinberg, , 2020; Elzayn et al., , 2019; Freund and Hssaine, , 2021), as we do. The algorithms developed in these papers, however, are non-adaptive: they decide on the entire allocation upfront, before observing any of the realized demand. In contrast, we consider a model where the decision-maker makes the allocation decision in each round after observing the number of arrivals. Freeman et al., (2017) considers a problem in which agents’ utilities are realized from an unknown distribution, and the budget resets in each round. They present algorithms for Nash social welfare maximization and discuss some of their properties. Our work is most closely related to (and indeed, builds upon) Sinclair et al., (2022), which first introduced the envy-freeness and efficiency tradeoff we are interested in. Our work considers the more challenging setting of perishable goods, which none of the aforementioned works consider.
Perishable resources.
@@
Though dynamic resource allocation of perishable goods has a long history in the operations research literature (see, e.g., Nahmias, (2011) for a comprehensive survey of earlier literature), to the best of our knowledge, the question of fairly allocating perishable goods has attracted relatively little attention. We highlight the few relevant papers below. Perry, (1999) and Hanukov et al., (2020) analyze FIFO-style policies for efficiency maximization in inventory models with Poisson demand and deterministic or Poisson perishing times. Motivated by the problem of electric vehicle charging, Gerding et al., (2019) consider an online scheduling problem where agents arrive and compete for a perishable resource that spoils at the end of every period, and as a result must be allocated at every time step. They consider a range of objectives, including: maximum total resource allocated, maximum number of satisfied agents, as well as envy-freeness. Bateni et al., (2022) similarly consider a setting wherein an arriving stream of goods perish immediately. Recent empirical work by Sengul Orgut and Lodree, (2023) considers the problem of a food bank equitably and efficiently allocating perishable goods under complete information. Their case study on data from a partnering food bank numerically validates our theoretical results: in low-budget settings, there is little or no benefit to increasing inequity to counteract the inefficiency due (in part) to spoilage. In contrast to these latter papers, the model we consider locates itself within the smaller category of inventory models in which products have random lifetimes. The majority of these assume that items have exponentially distributed or geometric lifetimes (Bakker et al., , 2012).

### v2

@@
[…]
MNSC-0001-2024.00
[…]
Banerjee, Hssaine, and Sinclair
[…]
Online Fair Allocation of Perishable Resources
[…]
Online Fair Allocation of Perishable Resources
[…]
Sid Banerjee
[…]
Chamsi Hssaine
[…]
Sean R. Sinclair
[…]
We consider a practically motivated variant of the canonical online fair allocation problem: a decision-maker has a budget of perishable resources to allocate over a fixed number of rounds. Each round sees a random number of arrivals, and the decision-maker must commit to an allocation for these individuals before moving on to the next round. The goal is to construct a sequence of allocations that is envy-free and efficient. Our work makes two important contributions toward this problem: we first derive strong lower bounds on the optimal envy-efficiency trade-off, demonstrating that a decision-maker is fundamentally limited in what she can hope to achieve relative to the no-perishing setting; we then design an algorithm achieving these lower bounds which takes as input $(i)$ a prediction of the perishing order, and $(ii)$ a desired bound on envy. Given the remaining budget in each period, the algorithm uses forecasts of future demand and perishing to adaptively choose from one of two carefully constructed guardrail quantities. We demonstrate our algorithm’s strong numerical performance — and state-of-the-art, perishing-agnostic algorithms’ inefficacy — on simulations calibrated to a real-world dataset.
[…]
Online Resource Allocation, Fairness, Perishability
@@ 1 Introduction
Despite a consistent decline in food insecurity in the United States over the past decade, 2022 saw a marked increase in individuals struggling to access enough food to fulfill basic needs. A recent report by the U.S. Department of Agriculture found that over 44 million individuals faced some form of hunger in 2022 — 45% more than the previous year (Godoy 2023). Due in part to rising food prices and the rolling back of pandemic-era social security measures, this disturbing statistic has further underscored the important role of local food banks; for instance, the Feeding America network of food banks, food agencies, and local food programs distributed over 5.2 billion meals that same year (Feeding America 2022).
@@
In distributing food throughout their operating horizon, food banks have two competing objectives: distributing as much food as possible to communities in need, and ensuring equitable access to donations. This tension has attracted much attention in the operations literature, with recent work characterizing the fundamental trade-offs between fairness and efficiency in sequential allocation problems (Bertsimas et al. 2011, Donahue and Kleinberg 2020, Lien et al. 2014, Manshadi et al. 2023, Sinclair et al. 2023, Onyeze et al. 2025b). Understanding such trade-offs in theory is useful, as they allow a system designer to recognize and choose their desired operating point, balancing the loss in efficiency and equity. Despite the useful insights derived in prior work, to the best of our knowledge an important reality of food bank operations remains overlooked: waste due to spoilage of perishable goods.
Perishable goods constitute a substantial portion of food donations. For instance, 35% of the food distributed annually by the Food Bank of Central and Eastern North Carolina is perishable (Sengul Orgut and Lodree 2023). Similarly, in 2024 alone, the Los Angeles Regional Food Bank distributed over 38 million pounds of fresh produce, representing 29% of the total amount of food distributed throughout the year (Los Angeles Regional Food Bank 2025). At the same time, the effective distribution of perishable goods has long been a significant challenge for these organizations. For example, in 2020 Feeding America reported that 3% of the food received by its partner agencies was disposed of due to spoilage. Similar percentages were reported by other food rescue organizations in the United States, with Boston- and Rochester-based organizations respectively reporting that 2.4% and 6% of all received food was wasted due to spoilage (U.S. Environmental Protection Agency 2020).
These figures highlight a persistent operational gap between the supply of perishable goods and their successful distribution. The reason for this gap is two-fold. First, food banks lack control over the kinds of donated items, meaning that the supply of goods that are about to perish can far exceed the demand for these goods. Ohls et al. (2002) highlighted this as a significant concern: in interviews with over 300 food banks, they found that around 20% of food banks distribute less than 90% of the food they receive, partly due to this issue. Second, even when this is not the case, true expiration dates are frequently unknown. Rather, product packaging commonly indicates “best-by,” “use-by,” or “sell-by” dates, after which products are still safe to consume (U.S. Department of Agriculture, Food Safety and Inspection Service 2019). So, while these dates may provide an informative signal as to when goods will perish (e.g., one week after the “best-by” date), the actual time at which this occurs is not known ex-ante. This problem is further exacerbated within the context of food banks, where individuals are encouraged to donate perishables that are past their “best-by” dates (City Harvest 2019, WQAD News 8 2025). Not only does this reduce the amount of time food banks have to allocate these items before they spoil, but it also creates significant uncertainty in the quality of the donations (Zou et al. 2025).
Intuitively, the equity-efficiency trade-off should be exacerbated in the presence of perishables: while equity requires a decision-maker to allocate conservatively to preserve resources for future demand (Sinclair et al. 2023), perishability starts a “race against time.” As goods perish due to a slow allocation rate, not only is efficiency harmed, but so may be equity, since spoilage runs the risk of a decision-maker running out of items, with nothing left to give out by the end of the operating horizon. Thus motivated, this paper seeks to answer the following questions:
@@
Do established equity-efficiency trade-offs in dynamic environments persist in the presence of perishable goods? If not, what limits do they impose on fair and efficient allocation? Can we design policies that perform well under these limits?
@@ 1.1 Our Contributions
We consider a model in which a decision-maker has a fixed budget of items (also referred to as goods, or resources) to be allocated over $T$ discrete rounds. An a priori unknown number of individuals arrives in each period, each seeking a share of goods. Each agent is characterized by an observable type (drawn from a known, potentially time-varying distribution), associated with a linear utility function over the received allocation. Moreover, each unit of good has a stochastic perishing time (similarly drawn from a known distribution, independent of incoming arrivals), after which the good spoils and can no longer be allocated. The decision-maker allocates goods according to a fixed ordering (also referred to as perishing prediction, or allocation schedule). For instance, if each good has an associated “best-by” date, in practice the decision-maker may allocate the goods in increasing order of these dates. The goal is to find a policy that trades off between three ex-post metrics:
@@
For this setting, we first characterize the fundamental limits that perishability places on any online algorithm. To see the considerable impact that perishability can have on counterfactual envy, consider an extreme scenario in which all items perish at the end of the first round. Clearly, there is no hope of achieving low envy in such a setting, since future demand can never be satisfied. Our first main contribution identifies a necessary and sufficient condition on the joint perishing and arrival distribution for low counterfactual envy to be an achievable desideratum (Theorem 3.3). From a managerial perspective, this characterization — which at a high level states that the cumulative perishing must lag behind cumulative arrivals — can be viewed as guidance on the composition of perishable donations that the decision-maker should accept. It moreover underscores the importance of leveraging joint information over perishing and demand: if demand is back-loaded, perishing times should be concentrated late in the horizon; if most demand arrives early, however, early perishing times are acceptable.
For this class of processes, which we term offset-expiring, zero spoilage occurs if goods are allocated in the exact order in which they perish. Due to stochasticity in goods’ perishing times, however, the decision-maker’s allocation schedule may be misaligned with the realized perishing times of goods. For instance, while one good may have an earlier “best-by” date than another, the latter good may spoil earlier. Alternatively, if the decision-maker allocates goods in increasing order of expected perishing time, the variance of the perishing times may cause goods with later expected perishing times to perish earlier than others. We show that the inherent aggressiveness of the perishing process (i.e., in cases where the process is not offset-expiring) and inaccuracies in the allocation schedule pose an insurmountable barrier to any algorithm’s performance by deriving lower bounds on any algorithm’s envy and inefficiency. These lower bounds are parameterized by an unavoidable loss due to spoilage, also referred to as the perishing-induced loss and denoted by $\mathcal{L}^{\textsf{perish}}$ (Definition 3.7), implicitly characterized as the solution to a fixed-point equation (Theorem 3.8). In contrast to the no-perishing setting, in which the only source of loss is exogenous uncertainty in demand, in our setting the loss due to spoilage is endogenous: it crucially depends on the rate at which the algorithm allocates items. This endogeneity poses significant challenges in the analysis of any adaptive algorithm; designing a tractable approach to analyzing ex-post spoilage is the main technical contribution of our work. Additionally, the lower bounds we derive give rise to the key insight that, contrary to the “race against time” intuition under which a decision-maker must increase the allocation rate to prevent avoidable spoilage, achieving low hindsight envy in the presence of unavoidable spoilage requires a decision-maker to potentially allocate significantly less than the proportional allocation. Hence, perishability throws a wrench into the well-studied envy-efficiency trade-off: while hindsight and counterfactual envy are aligned in the no-perishing setting, these two may be at odds when goods spoil, since only high counterfactual-envy solutions may yield low hindsight-envy. The tension between efficiency and equity is exacerbated for the same reason, relative to the classical setting.
In our final technical contribution, we leverage these insights to construct an adaptive threshold algorithm (Algorithm 1) that achieves these lower bounds (Theorem 4.1). Our algorithm takes as input $(i)$ the fixed allocation schedule, $(ii)$ a desired upper bound on hindsight envy $L_{T}$, and $(iii)$ a high-probability parameter $\delta$. Given these inputs, it computes a high-probability lower bound on a budget-respecting zero-hindsight-envy solution, and an “aggressive” efficiency-improving allocation that is $L_{T}$ away. In each round the algorithm chooses which of the two quantities to allocate to each individual, cautiously doing so by constructing pessimistic forecasts of future arrivals and perishing. While this algorithm is similar in flavor to state-of-the-art algorithms for the no-perishing setting (Sinclair et al. 2023), the main challenge it contends with is forecasting endogenous future spoilage. We overcome this challenge by leveraging an exogenous “slow” consumption process that tractably decouples past allocations from future perishing. Crucially, we show that future spoilage under this slow process upper bounds our algorithm’s future spoilage.
@@
Figure 1: Graphical representation of Theorems 4.1 and 3.8, illustrating the envy-efficiency trade-off ($\Delta_{\text{\it efficiency}}$ vs. $\Delta_{\text{\it EF}}$) achieved by Perishing-Guardrail (Algorithm 1) in the “high-envy, low-perishing” regime. Here, $\mathcal{L}^{\textsf{perish}}$ is used to denoted the unavoidable perishing-induced loss (Definition 3.7). The vertical and horizontal dotted lines represent the impossibility results due to demand and perishing uncertainty, respectively. The region below the solid line represents the impossibility due to the envy-efficiency trade-off; the green region is the achievable region for Perishing-Guardrail.
Our bounds give rise to three salient regimes that dictate the extent to which a decision-maker is restricted in leveraging inequity to improve efficiency (and vice versa). These three regimes are a function of both $L_{T}$ and the unavoidable perishing-induced loss dictating our lower bounds, and at a high level are as follows:
@@
    Low-envy: There are no efficiency gains from deviating from the equitable solution, where all individuals receive the same amount.
@@
    High-envy, high-perishing: Similarly here, inefficiency is invariant to $L_{T}$.
@@
    High-envy, low-perishing: Inefficiency decreases as $1/L_{T}$, until reaching the unavoidable perishing-induced loss throughout the horizon, $T\mathcal{L}^{\textsf{perish}}$. This regime is represented pictorially in Figure 1.
We conclude our theoretical results by providing explicit characterizations of the unavoidable perishing-induced loss for important special cases (Propositions 4.16, 4.17, 4.18 and 4.19). These characterizations shed light on the salient drivers of this loss: $(i)$ how aggressively goods perish relative to the rate at which they are demanded, and $(ii)$ whether the allocation schedule approximately preserves the order induced by the true perishing times. Our results reveal that the variability of the perishing distribution plays a crucial role in both cases.
We complement our theoretical bounds by testing the practical performance of our algorithm on both synthetic and real-world datasets. Our experiments show that the unfairness required to achieve efficiency gains is order-wise larger than in settings without perishable resources. Additionally, they underscore the weakness of perishing-agnostic online algorithms. We observe that these latter algorithms are incapable of leveraging unfairness to improve efficiency across a variety of perishing regimes. In contrast to these, our algorithm’s construction of a perishing-aware baseline allocation is necessary to mitigate stockouts across all — rather than simply worst-case — instances. These include instances where offset-expiry fails to hold with high probability, as is the case in the real-world dataset we use to calibrate our experiments (Keskin et al. 2022). Perhaps most surprisingly, despite our baseline allocation being significantly lower than that of algorithms that don’t take into account perishability, our algorithm is more efficient than these more aggressive algorithms, in addition to being more fair. This observation contradicts the “race against time” intuition that aggressive allocations are necessarily more efficient than cautious ones. Finally, we numerically explore the question of selecting practical allocation schedules, with the goal of minimizing the unavoidable loss due to perishing. Our main managerial insight is that ordering items according to a high-probability lower bound on their perishing time is a robust, interpretable choice for decision-makers; it naturally takes into account items’ expected perishing times, all the while hedging against the variability of the perishing process.
Finally, we note that while our work is motivated by food bank operations, perishability is also an important consideration in medical applications, such as blood bank operations (Bar-Lev et al. 2017) and vaccine distribution (Manshadi et al. 2023). For these applications, our modeling framework is applicable to settings in which allocation occurs in rural or remote areas, where uncertainty around storage conditions (e.g., temperature and humidity) may affect the lifetime of these goods.
@@
We next survey the related literature. We present the model in Section 2, and formalize the limits of perishability in Section 3. In Section 4 we design and analyze an algorithm that achieves the derived lower bounds, illustrating the impact of the quality of the allocation schedule and the aggressiveness of the perishing process for important special cases in Section 4.4. We conclude with numerical experiments in Section 5.
@@ 1.2 Related Work
Fairness in resource allocation has a long history in the economics and computation literature, beginning with Varian’s seminal work (Varian 1974, 1976). More recently, there has been work studying fairness in operations, including resource-level fairness through load balancing (Chen et al. 2022, Hssaine et al. 2024, Balseiro et al. 2025, Freund et al. 2025), pricing (Cohen et al. 2022, den Boer et al. 2022), incentive design (Freund and Hssaine 2025, Hssaine et al. 2025), algorithmic hiring and online selection (Salem and Gupta 2024, Hu et al. 2025), and societal systems more generally (Gupta and Kamble 2021, Lyu et al. 2025). We highlight the most closely related works below, especially as they relate to online fair allocation; see Aleksandrov and Walsh (2019b) for a survey.
@@
There exists a long line of work in which non-perishable resources become available to the decision-maker online, whereas agents are fixed (Benade et al. 2018, Aleksandrov et al. 2015, Mattei et al. 2017, 2018, Aleksandrov and Walsh 2019a, Banerjee et al. 2022, Bansal et al. 2020, Bogomolnaia et al. 2022, He et al. 2019, Aziz et al. 2016, Zeng and Psomas 2020). These models lie in contrast to the one we consider, wherein resources are fixed and individuals arrive online. Papers that consider this latter setting include Kalinowski et al. (2013), which studies welfare maximization with indivisible goods; Gerding et al. (2019) considers a scheduling setting wherein agents have fixed and known arrival and departure times, as well as demand for the resource; Hassanzadeh et al. (2023) allows individuals to arrive in multiple timesteps. A series of papers also consider the problem of fair allocation with minimal disruptions relative to previous allocations, as measured by a fairness ratio, a competitive ratio analog of counterfactual envy in our setting (Friedman et al. 2017, Cole et al. 2013, Friedman et al. 2015). Other works design algorithms with attractive competitive ratios with respect to Nash Social Welfare (Azar et al. 2010, Banerjee et al. 2022) and the max-min objective (Lien et al. 2014, Manshadi et al. 2023). Ma et al. (2025) recently showed how forward-backwards contention resolution schemes can be leveraged to obtain competitive ratios for online rationing in non-perishable settings.
@@
The above papers situate themselves within the adversarial, or worst-case, tradition. A separate line of work considers fair resource allocation in stochastic settings (Donahue and Kleinberg 2020, Elzayn et al. 2019), as we do. The algorithms developed in these papers, however, are non-adaptive: they decide on the entire allocation upfront, before observing any of the realized demand. In contrast, we consider a model where the decision-maker makes the allocation decision in each round after observing the number of arrivals. Freeman et al. (2017) considers a problem in which agents’ utilities are realized from an unknown distribution, and the budget resets in each round. They present algorithms for Nash social welfare maximization and discuss some of their properties. Our work is most closely related to (and indeed, builds upon) Sinclair et al. (2023), which first introduced the envy-efficiency tradeoff we are interested in. Our work considers the more challenging setting of perishable goods, which none of the aforementioned works consider.
Finally, we note the existence of a rich line of work that takes a mechanism design approach to fair division, dating back to the 1940s (Steinhaus 1948). We refer the reader to Moulin (2004), Chapters 11-13 of Brandt et al. (2016), and Chapter 11 of Kroer (2026) for textbook treatments of the topic. Generally speaking, the main objective of these works is the design of nonmonetary mechanisms that fairly allocate a set of resources amongst strategic agents with private preferences for these goods. Much of the work has focused on one-shot (i.e., static) settings, for which there exist strong impossibility results precluding the existence of fair, efficient, and incentive compatible mechanisms (Budish 2011). Notable works that derive approximately fair and incentive-compatible mechanisms for these settings include Budish (2011) and Brânzei et al. (2017). In dynamic settings, models of fair allocation under incentives involve the same set of agents competing for resources across time, which differs from the arrival model we consider here. Recent works that consider this setting include Onyeze et al. (2025a), who study equilibrium existence, and Fikioris et al. (2024), who propose a generalization of the popular dynamic max-min fair allocation mechanism to achieve strong fairness and approximate incentive compatibility guarantees.
We conclude by highlighting the closely related work of Prendergast (2022), which empirically studies the virtual currency-based auction mechanism used by Feeding America to allocate food to its members. This work finds that, compared to a previous queuing mechanism under which food banks were being allocated an equal amount of random food through time, the auction mechanism allows food banks to express preferences across different types of food. As it relates to our work, it would be interesting and practically relevant to study how $(i)$ the timing of allocations, and $(ii)$ varying preferences across resources induce strategic behavior. We leave this for future work.
Dynamic allocation of perishable resources.
@@
Though dynamic resource allocation of perishable goods has a long history in the operations research literature (see, e.g., Nahmias (2011) for a comprehensive survey of earlier literature), to the best of our knowledge, the question of fairly allocating perishable goods has attracted relatively little attention. We highlight the few relevant papers below. Perry (1999) and Hanukov et al. (2020) analyze FIFO-style policies for efficiency maximization in inventory models with Poisson demand and deterministic or Poisson perishing times. Motivated by the problem of electric vehicle charging, Gerding et al. (2019) consider an online scheduling problem where agents arrive and compete for a perishable resource that spoils at the end of every period, and as a result must be allocated at every time step. They consider a range of objectives, including: maximum amount of resources allocated, maximum number of satisfied agents, as well as envy-freeness. Bateni et al. (2022) similarly consider a setting wherein an arriving stream of goods perish immediately. Recent empirical work by Sengul Orgut and Lodree (2023) considers the problem of a food bank equitably and efficiently allocating perishable goods under complete information. Their case study on data from a partnering food bank numerically validates our theoretical results: in low-budget settings, there is little or no benefit to increasing inequity to counteract the inefficiency due (in part) to spoilage. In contrast to these latter papers, the model we consider locates itself within the smaller category of inventory models in which products have random lifetimes. The majority of these assume that items have exponentially distributed or geometric lifetimes (Bakker et al. 2012).